A 4.0 kg bundle starts up a 30° incline with 200 J of kinetic energy. How far will it


slide up the incline if the coefficient of kinetic friction between the bundle and the incline is


0.42?

1
Expert's answer
2021-12-16T11:33:19-0500

Apply conservation of energy. The initial energy will be spent to lift the block up the slope (increase its potential energy) and overcome friction along the incline:


"E=mgL\sin\theta+\mu mgL\cos\theta,\\\space\\
L=\frac E{mg(\sin\theta+\mu \cos\theta)}=5.9\text{ m}."